On heating water, bubbles being formed at the bottom of the vessel detach and rise. Take the bubbles to be spheres of radius R and making a circular contact of radius r with the bottom of the vessel. If r

Grade:10

2 Answers

Arun
4 years ago
Dear Anmol
 
Question is incomplete. Please check and repost this. You can also attach an image.
I will be happy to help you.
Khimraj
4 years ago
(2πrT)sin⁡θ=(4/3)πR3ρwg
T×rR×2πr=(4/3)πR3ρwg
r2=(2/3)R4ρwg/T
r=R2wg/3T)1/2 is correct.
Hope it clears..........................................................